# Predicting the single-site and multi-site event discrimination power of dual-phase time projection chambers

### v2

@@ 2.3 TPC resolution predictions
The predicted resolution is reported between 20 and 1000,000 detected photons, with the lower bound corresponding to the typical number of photons detected in a single electron event in a dual-phase xenon TPC and the upper bound approximating the size of 2.5 MeV gamma ray or electron events (Sec. 4.1). The resolution is observed to improve with increased photon counts in the form of $1/\sqrt{m_{1}}$, which is expected from Poisson distributions. With a 2" photosensor size, at 20 detected photons the predicted position resolution is $\sim$1 cm, which is a factor of 2 better than that demonstrated by the LUX experiment using 2" PMTs at a similar photon number [14]. When the LUX LRF function is used in a similar theoretical study (Appendix B), the predicted SS position reconstruction resolution is within 20% of experimental observations. The excellent agreement between this study and the LUX experimental result provides a strong confirmation that FI can be applied to position reconstruction in dual-phase xenon TPCs.
Figure 3 (right) shows the predicted $x$ position resolution for different photosensor sizes, from 3" diameter in the top to 6 mm in the bottom. Here we only plot the results for event position A which has the worst sensitivity. As the sensor size decreases, the predicted position resolution improves approximately linearly until the sensor size becomes smaller than the light source-sensor separation. This is again the inverse square law dependence of the photon intensity for isotropic light emission, which dominates the light spread once the light source is further away from the photosensor plane than the sensor pixel size. With the TPC model described in Sec. 2.1, the SS reconstruction capability is comparable for 15 mm and 6 mm sensor pixel sizes. However, a more significant improvement can be achieved for detectors that optimize the light source-sensor separation for smaller sensor pixels, to be discussed in Sec. 4.2.
@@
To study the dependence of $\sigma(d)$ on both $d$ and the number of photons at the two vertices, we compute the average $\sigma(d)$ over all $\phi$ angles, and only choose position A for the first vertex to represent the system performance. Position A is chosen for its conservative performance and its azimuthal symmetry in the placement of the second vertex. Figure 5 shows the resulting $\sigma(d)$ as a function of $d$ for 2" (left) and 6 mm sensors (right). We assume a total of 500 photons detected split between the two vertices with ratios of 1-1 (red solid line), 1-2 (green dotted line), 1-4 (blue dashed line), and 1-9 (magenta dot-dashed line). As expected, $\sigma(d)$improves with increasing $d$. The fine features observed in the case of 2" photosensors are a result of the second vertex crossing the boundaries between neighboring sensor pixels. For 6 mm studies, $\sigma(d)$ decreases monotonically with increasing $d$ without any oscillating features because the photosensor pixelation is a subdominant effect due to the large distance between the light source and the sensors. At a fixed total photon count, large disparities in vertex brightness degrade the $d$ reconstruction, as the hit pattern produced by the dimmer vertex can be shadowed by that of the brighter one, causing its position information to be partially lost.
@@
Figure 5: Predicted $\sigma(d)$ as a function of vertex separation ($d$), with 500 total detected photons for MS2 events with 2" (left) and 6 mm (right) photosensor sizes. The lines indicate different photon budget splitting between the two vertices: 1-1 splitting (solid line), 1-2 splitting (dotted), 1-4 splitting (dashed), 1-9 splitting (dot dashed). The red line represents $\sigma(d)=d$. The diverging trend near $d=0$ reflects an instability of the FI matrix and is unphysical.
@@
For very small $d$ values, FI predicts $\sigma(d)$ values in the range of tens of cm, which are even larger than the spread of photon pattern for SS events in our TPC model. This prediction is unphysical and represents an instability of the FI matrix in near-degeneracy conditions. As discussed in Sec. 2.2, when two vertices are close to each other, a degeneracy between parameters describing the vertices emerges; changes in a parameter for one vertex may be compensated by a proportionate change in parameters for the other vertex. Mathematically, degeneracies of parameters cause the FI matrix which is constructed from likelihood score covariance, to become ill-conditioned. With a complete degeneracy between two parameters, the FI matrix is non-invertible and the CRLB limits would diverge. For these unphysical $\sigma(d)$ values, the condition number of the FI matrix, calculated as the ratio of the largest eigenvalue to the smallest, is observed to increase by orders of magnitude even with a millimeter-level change in $d$. In this work, we discard values of $\sigma(d)$ larger than 10cm. The validity of this approach will be further studied with the simulation and reconstruction efforts presented in Sec. 3.
@@
We define a benchmark MS2 resolution $\delta_{d}$ as the value $d$ at which $\sigma(d)=d$, i.e., when the precision to reconstruct the distance between the two vertices is equal to the distance itself. In this convention, at $d=\delta_{d}$, there is a $\sim$16% probability for an unbiased estimator to reconstruct the $d$ value to be negative, while an unbiased estimator is expected to reconstruct a mean $d$ value of 0 for SS events  [^2] Our FI calculation does not consider bounds for parameters (unbiased), so the inter-vertex distance $d$ (and the photon count $m$) may take on both positive and negative values. . This corresponds to the scenario of 16% MS leakage with 50% SS acceptance with a cut of $d<0$ and will be used as the benchmark for comparison with reconstruction results in Sec. 3. In Fig. 5, the benchmark resolution values are obtained at the intercepts of the $y=x$ line with the $\sigma(d)-d$ curves. For 2"-size sensors $\delta_{d}$=1.2 cm when 500 photons are detected and the photons are equally distributed between the two vertices. $\delta_{d}$ increases to 3.1 cm when one vertex is $\sim$10% as bright as the other, even when the same number of total photons is detected. Similar observations are made for the 6 mm sensor pixel sizes, except that both $\sigma(d)$ and its dependence on sensor layout is reduced thanks to the finer pixelation of the photosensor array.
@@
Figure 6: Benchmark $d$ reconstruction resolution as a function of number of photons detected for 1-1 (left) and 1-9 (right) photon budget splitting between the two vertices. Different lines indicate different photosensor sizes - from top to bottom: 3", 2", 1", 15 mm, and 6 mm circular photosensors.
@@
Figure 6 shows the benchmark MS2 resolution $\delta_{d}$ calculated as a function of the number of photons detected, for different photosensor pixel sizes: 3", 2", 1", 15 mm, and 6 mm (from top to bottom); the figure on the left assumes a 1-1 photon splitting between the two vertices while the figure on the right assumes a 1-9 photon splitting. Generally, collecting more photons and using smaller sensor pixels reduce uncertainties in the estimated separation between the two vertices, but the improvement diminishes when the sensor size becomes smaller than the distance between the light source and the photosensor array. This is consistent with the results for the SS event reconstruction predictions. In addition, vertices with comparable brightness are easier to resolve than those with large disparities; a deterioration factor of 3 is observed for the 1-9 photon splitting scenario compared to the equal photon splitting scenario.
@@
Lastly, we point out that the smallest position separation between two vertices that can be resolved is significantly larger than the predicted position resolution for SS events, when all considered vertices have the same photon number. For example, in Fig. 3, with 250 photons detected, a SS event can be reconstructed with 3 mm accuracy for 2" photosensors, but when two vertices, each with detecting 250 photons, are present, the best reconstructed resolution on the separation between the two is 1.2 cm. This factor of 4 discrepancy is much larger than that of $\sqrt{2}$ expected by subtracting one vertex position from the other. This results from the ambiguity in associating a detected photon with a certain vertex. When there is only one vertex in the system, each detected photon carries definitive information about its origin, whereas for multiple vertices hit patterns significantly overlap and photons from both vertices may have similar probability to hit the same sensor pixel. This contributes to the uncertainty in locating the position for all overlapping vertices.
@@
To clearly demonstrate this point, we study the position reconstruction uncertainty for a vertex with and without a second vertex nearby. Fig. 7 shows the theoretically predicted position resolution for a vertex with 100 detected photons as a function of the distance at which a second vertex of equal (dashed line) and 9 times (solid line) photon count is placed. When the two vertices are far apart, the predicted position resolution approaches the SS prediction (dotted line), but when the two vertices are close to each other, the best reconstruction accuracy deviates substantially from the SS limit.
3 Reconstruction and discrimination methods
@@
In this section, we test position reconstruction and SS/MS discrimination by implementing methods including centroiding, maximum likelihood estimator (MLE), and convolutional neural network (CNN). The results obtained will be compared to theoretical predictions. The studies are performed with a simplified detector model employing 2" PMTs but the same method can be applied to other detector models described in Sec. 2 or more realistic detectors provided their light responses are known from simulation or direct measurements.
3.1 Centroiding reconstruction
@@
For a given sensor hit pattern, the mean position of all recorded photon hits provides a straightforward estimate of the event location. This method is often referred to as centroiding or center of mass. Fig. 8 (left) shows the $x_{1}$ centroid position for SS events as a function of the true $x_{1}$ position for 100 (grey points) and 10,000 (black points) detected photons, using a 2" sensor pixel size. Generally, the centroid position closely tracks the real event position, and the precision improves with increasing number of photons detected. However, even with 10,000 detected photons, the centroid position may still deviate from the true event position with a pattern corresponding to the sensor layout. This discrepancy originates from the inaccurate photon detection position estimated using discrete pixel center positions and can be mitigated by reducing the pixel size or increasing the light source-sensor separation. For actual detectors with a finite sensor array size, the centroid position will carry an additional bias from the asymmetric detection of photons, which usually pulls the reconstructed event position toward the center of the detector.
Figure 8: Left - centroid position reconstruction for SS events with 100 (gray) and 10,000 (black) photons detected using 2" sensors; Right - simulated photon spread in the sensor plane for a MS2 event, where the asterisks mark the vertex positions and the photons from different vertices are indicated with different colors; the arrows correspond to the principal and secondary orientations.
In addition to being a biased reconstruction algorithm, centroiding, by definition, provides only one position estimate and cannot reconstruct MS event locations. However, MS2 events could produce a photon hit pattern with larger variance along the direction connecting the two vertices (hereafter referred to as the principal direction) compared to its orthogonal direction, as illustrated in Fig. 8 (right). To quantify this asymmetric variance for SS and MS events, we use the Principal Component Analysis (PCA) to identify the principal axis direction. Mathematically, PCA diagonalizes the covariance matrix and rotates the coordinates to align with the principal direction. In addition to computing the rotation angle, PCA also calculates the eigenvalues of the covariance matrix, which correspond to the variance in photon spread in the principal and secondary directions. We will refer to the square roots of these variance values as the principal and secondary spreads.
Figure 9 (left) shows the distributions of principal (blue) and secondary (black) spread for double vertices separated by 0, 3cm and 6cm with a total of 500 detected photons. Here we assume a sensor pixel size of 2" and an 1-1 photon splitting between the two vertices. Due to statistical fluctuations, the PCA algorithm proposes principal and secondary directions even for SS events (0 separation between vertices). Therefore, the appearance of asymmetric variance, by itself, cannot be used to discriminate SS and MS events. Instead, we use the displacement of the mean principal spread from the SS mean to estimate how well MS events may be distinguished from SS events. This method takes statistical fluctuations of SS variance distribution into consideration. When this displacement equals the width of the principal spread distribution, we obtain a 50% SS acceptance while allowing 16% MS contamination. The corresponding vertex separation can be directly compared to the benchmark MS2 resolution $\delta_{d}$ defined in Sec. 2.3. The PCA-derived $\delta_{d}$ values are compared to that predicted by FI, as a function of detected photon count, in Figure 9 (right). The PCA-derived resolution values follow the same trend as the theoretical predictions, but are approximately twice larger at the same photon statistics.
3.2 Maximum likelihood reconstruction
@@
As demonstrated by the ZEPLIN-III and LUX experiments [13, 14], MLE can accurately reconstruct the position for SS events using a sensor’s response to light sources of known locations. With an array of 2" PMTs, LUX achieved a position resolution of 2.1 cm for single electron pulses with $\sim$20 detected photons, or $\sim$20% from our theoretical prediction in Appendix B. In this section, we use a similar MLE to reconstruct the positions for SS events and to estimate the inter-vertex separation for MS2 events. Given a sensor light response function of $\eta(r)$, the likelihood function for a given measurement can be calculated as
@@
Figure 10: Left - Maximum likelihood reconstruction for SS events with 100 and 10000 photons detected using 2" sensor pixels. Right - reconstructed position resolution for SS events right under a PMT (position A) as a function of photons detected.
@@
Figure 10 (left) shows the MLE-reconstructed $x_{1}$ position for SS events of 100 and 10,000 detected photons for a 2" photosensor size. Compared to the centroiding method shown in Fig. 8, the reconstruction uncertainty is signifiantly reduced, and more importantly, the reconstruction does not exhibit any significant deviation from the true event position. This is more evident for the result obtained with 10,000 detected photons which has much reduced statistical fluctuation. The figure in the right shows the resolution of reconstruction positions as a function of detected photons (solid black), in comparison to that predicted by FI (dashed red), the two of which are nearly identical. Because FI and CRLB are based on likelihood and the MLE reconstruction maximizes the likelihood, it is not surprising that the theoretical limit can be achieved for a relatively simple SS system when there is no additional uncertainty. A similar study for the LUX detector (Appendix B) predicts $\sim$20% better SS resolutions than experimental observations; the discrepancy may be attributed to the omission of experimental effects such as reflection off detector walls and non-Poissonian photon fluctuations in the idealized LRF.
For MS2 events, we use the PCA-estimated principal direction and variance to initialize the MLE fit. For well-separated MS2 events, the MLE is able to correctly reconstruct the vertex positions and light intensities. However, as the vertices move closer to each other, the MLE fit becomes unstable. Instead of examining the best-fit $d$ values and their estimated uncertainties, we quantify the goodness of MLE fits for simulated MS2 events with the likelihood ratio, which is defined as
$q=-2\textrm{log}\left(\frac{\mathcal{L}(\hat{m}_{1},\hat{x}_{1},\hat{y}_{1},m_{2}=0|n_{k})}{\mathcal{L}(\hat{m}_{1},\hat{x}_{1},\hat{y}_{1},\hat{m}_{2},\hat{d},\hat{\phi}|n_{k})}\right)$ (3.2)

### v3

@@ 2.3 Position resolution predictions
The predicted resolution is reported between 20 and 1,000,000 detected photons, with the lower bound corresponding to the typical number of photons detected in a single electron event in a dual-phase xenon TPC and the upper bound approximating the size of 2.5 MeV gamma ray or electron events (Sec. 4.1). The resolution is observed to improve with increased photon counts in the form of $1/\sqrt{m_{1}}$, which is expected from Poisson distributions. With a 2" photosensor size, at 20 detected photons the predicted position resolution is $\sim$1 cm, which is a factor of 2 better than that demonstrated by the LUX experiment using 2" PMTs at a similar photon number [13]. When the LUX LRF function is used in a similar theoretical study (Sec. 3.1.1), the predicted SS position reconstruction resolution is within 20% of experimental observations. The excellent agreement between this study and the LUX experimental result provides a strong confirmation that FI can be applied to position reconstruction in dual-phase xenon TPCs.
Figure 3 (right) shows the predicted $x$ position resolution for different photosensor sizes, from 3" to 6 mm in diameter. Here we only plot the results for event position A which has the worst sensitivity. As the sensor size decreases, the predicted position resolution improves approximately linearly until the sensor size becomes smaller than the light source-sensor separation. This is again the inverse square law dependence of the photon intensity for isotropic light emission, which dominates the light spread once the light source is further away from the photosensor plane than the sensor pixel size. With the TPC model described in Sec. 2.1, the SS reconstruction capability is comparable for 15 mm and 6 mm sensor pixel sizes. However, a more significant improvement can be achieved for detectors that optimize the light source-sensor separation for smaller sensor pixels, to be discussed in Sec. 4.2.
@@
To study the dependence of $\sigma(d)$ on both $d$ and the number of photons at the two vertices, we compute the average $\sigma(d)$ over all $\phi$ angles, and only choose position A for the first vertex to represent the system performance. Position A is chosen for its conservative performance and its azimuthal symmetry in the placement of the second vertex. Figure 5 shows the resulting $\sigma(d)$ as a function of $d$ for 2" (left) and 6 mm sensors (right). We assume a total of 500 photons detected split between the two vertices with ratios of 1:1 (solid black), 1:2 (dotted black), 1:4 (dashed black), and 1:9 (dot-dashed black). As expected, $\sigma(d)$ improves with increasing $d$. The fine features observed in the case of 2" photosensors are a result of the second vertex crossing the boundaries between neighboring sensor pixels. For 6 mm studies, $\sigma(d)$ decreases monotonically with increasing $d$ without any oscillating features because the photosensor pixelation is a subdominant effect due to the large distance between the light source and the sensors. At a fixed total photon count, large disparities in vertex brightness degrade the $d$ reconstruction, as the hit pattern produced by the dimmer vertex can be shadowed by that of the brighter one, causing its position information to be partially lost.
@@
Figure 5: Predicted $\sigma(d)$ as a function of vertex separation ($d$), with 500 total detected photons for MS2 events with 2" (left) and 6 mm (right) photosensor sizes. The lines indicate different photon budget splitting between the two vertices: 1:1 splitting (solid line), 1:2 splitting (dotted), 1:4 splitting (dashed), 1:9 splitting (dot dashed). The red line represents $\sigma(d)=d$. The diverging trend near $d=0$ reflects an instability of the FI matrix and is unphysical.
@@
For very small $d$ values, FI predicts $\sigma(d)$ values in the range of tens of cm, which are even larger than the spread of photon patterns for SS events in our TPC model. This prediction is unphysical and represents an instability of the FI matrix in near-degeneracy conditions. As discussed in Sec. 2.2, when two vertices are close to each other, a degeneracy between parameters describing the vertices emerges; changes in a parameter for one vertex may be compensated by a proportionate change in parameters for the other vertex. Mathematically, degeneracies of parameters cause the FI matrix, which is constructed from likelihood score covariance, to become ill-conditioned. With a complete degeneracy between two parameters, the FI matrix has eigenvalues containing 0, and is thus non-invertible; as a result, the CRLB limits would diverge. For these unphysical $\sigma(d)$ values, the condition number of the FI matrix, calculated as the ratio of the largest eigenvalue to the smallest, is observed to increase by orders of magnitude even with a millimeter-level change in $d$. In this work, we exclude values of $\sigma(d)$ greater than 10 cm. The validity of this approach will be further studied with the simulation and reconstruction efforts presented in Sec. 3.
@@
We define a benchmark MS2 resolution $\delta_{d}$ as the value $d$ at which $\sigma(d)=d$, i.e., when the precision to reconstruct the distance between the two vertices is equal to the distance itself. At this distance, there is a $\sim$16% probability for an unbiased estimator to reconstruct $d$ as negative  [^2] Our FI calculation does not consider bounds for parameters (unbiased), so the inter-vertex distance $d$ (and the photon count $m$) may take on both positive and negative values. if we assume a Gaussian distribution for the obtained $d$ value. The same estimator would also predict a mean $d$ value of 0 for SS events. Therefore, a cut of $d<0$ would accept 50% SS events while allowing for a 16% MS leakage, which will be used as the benchmark for comparison with reconstruction results in Sec. 3. In Fig. 5, the benchmark resolution values are obtained at the intercepts of the $y=x$ line with the $\sigma(d)-d$ curves. For 2"-size sensors $\delta_{d}$=1.2 cm when 500 photons are detected and the photons are equally distributed between the two vertices. $\delta_{d}$ increases to 3.1 cm when one vertex is $\sim$10% as bright as the other, even when the same number of total photons is detected. Similar observations are made for the 6 mm sensor pixel sizes, except that both $\sigma(d)$ and its dependence on sensor layout is reduced thanks to the finer pixelation of the photosensor array.
@@
Figure 6: Benchmark $d$ reconstruction resolution as a function of number of photons detected for 1:1 (left) and 1:9 (right) photon budget splitting between the two vertices. Different lines indicate different photosensor sizes - from top to bottom: 3", 2", 1", 15 mm, and 6 mm circular photosensors.
@@
Figure 6 shows the benchmark MS2 resolution $\delta_{d}$ calculated as a function of the number of photons detected, for different photosensor pixel sizes: 3", 2", 1", 15 mm, and 6 mm (from top to bottom); the figure on the left assumes a 1:1 photon splitting between the two vertices while the figure on the right assumes a 1:9 photon splitting. Generally, collecting more photons and using smaller sensor pixels reduce uncertainties in the estimated separation between the two vertices, but the improvement diminishes when the sensor size becomes smaller than the distance between the light source and the photosensor array. This is consistent with the results for the SS event reconstruction predictions. In addition, vertices with comparable brightness are easier to resolve than those with large disparities; a deterioration factor of 3 is observed for the 1:9 photon splitting scenario compared to the equal photon splitting scenario.
@@
Lastly, we point out that the smallest position separation between two vertices that can be resolved is significantly larger than the predicted position resolution for SS events, when all considered vertices have the same photon number. For example, in Fig. 3, with 250 photons detected, a SS event can be reconstructed with 3 mm accuracy for 2" photosensors, but when two vertices, each with detecting 250 photons, are present, the benchmark resolution of the separation between the two is 1.2 cm. This factor of 4 discrepancy is much larger than that of $\sqrt{2}$ expected by subtracting one vertex position from the other. This results from the ambiguity in associating a detected photon with a certain vertex. When there is only one vertex in the system, each detected photon carries definitive information about its origin, whereas for multiple vertices hit patterns significantly overlap and photons from both vertices may have similar probability to hit the same sensor pixel. This contributes to the uncertainty in locating the position for all overlapping vertices.
@@
To clearly demonstrate this point, we study the position reconstruction uncertainty for a vertex with and without a second vertex nearby. Fig. 7 shows the theoretically predicted position resolution for a vertex with 250 detected photons as a function of the distance at which a second vertex of equal (dashed line) and 9 times (solid line) photon count is placed. When the two vertices are far apart, the predicted position resolution approaches the SS prediction (dotted line), but when the two vertices are close to each other, the best reconstruction accuracy deviates substantially from the SS limit.
3 Reconstruction and validation
@@
In this section, we test position reconstruction and SS/MS discrimination with our simple TPC model. Results obtained from methods such as centroiding, principal component analysis (PCA), maximum likelihood estimator (MLE), and convolutional neural network (CNN) are compared to theoretical predictions in Sec. 2.3 and available experimental results. All reconstruction studies are performed with 2" PMTs in the TPC model, and for MS events we only consider MS2 events with equal photon split between the two vertices. However, the same methods can be applied to other detector models studied in Sec. 2, or more realistic detectors provided their light responses are known from simulation or direct measurements.
3.1 Single-site position reconstruction
@@
For a given sensor hit pattern, the mean position of all recorded photon hits provides a straightforward estimate of the event location. This method is often referred to as centroiding or center of mass. Fig. 8 (left) shows the $x_{1}$ centroid position for SS events as a function of the true $x_{1}$ position for 100 (grey points) and 10,000 (black points) detected photons, using a 2" sensor pixel size. True $x_{1}$ positions are randomly sampled between -10 cm and 10 cm while keeping $y_{1}$ at 0. The sensor hit patterns are then generated using a toy Monte Carlo simulation that samples the parameterized LRF. Generally, the obtained centroid position closely tracks the true event position, and the precision improves with increasing number of photons detected. However, even with 10,000 detected photons, the centroid position may still deviate from the true event position with a pattern corresponding to the sensor layout. This discrepancy originates from the inaccurate photon detection position estimated using discrete pixel center positions and can be partially mitigated by reducing the pixel size or increasing the light source-sensor separation. For actual detectors with a finite size of sensor array, the centroid position will carry an additional bias from the asymmetric detection of photons, which usually pulls the reconstructed event position toward the center of the detector.
Figure 8: Left - centroid $x_{1}$ position reconstructed for SS events ($y_{1}$=0) with 100 (gray) and 10,000 (black) photons detected in the TPC model using 2" sensors; Right - the same as left but with the maximum likelihood reconstruction. The coordinates are defined in Fig. 1.
@@
The aforementioned bias with centroiding estimation is addressed in more sophisticated methods such as MLE. As demonstrated by the ZEPLIN-III and LUX experiments [19, 13], MLE can accurately reconstruct the position of SS events using a sensor’s response to light sources of known locations. Given a sensor response function of $\eta(r)$, the likelihood function for a given measurement can be calculated as
@@
Figure 9: Reconstructed position resolution for SS events right under a PMT (position A) in the TPC model with 2" PMTs as a function of photons detected.
@@
Figure 8 (right) shows the MLE-reconstructed $x_{1}$ position for SS events of 100 and 10,000 detected photons for 2" photosensors. Compared to the centroiding result in the left subfigure, the MLE reconstruction uncertainty is significantly reduced, and more importantly, the reconstruction does not exhibit any significant deviation from the true event position. This is more evident for the result obtained with 10,000 detected photons which has much reduced statistical fluctuation. Figure 9 shows the resolution of MLE reconstruction positions as a function of detected photons (solid black), in comparison to that predicted by FI (dashed red), the two of which are nearly identical. Because both FI and CRLB are based on likelihood and the MLE reconstruction maximizes the likelihood, it is not surprising that the theoretical limit can be achieved for a relatively simple SS system when there is no additional uncertainty. In an actual detector where the LRF is not exactly known or its system dependence is not fully captured by the model, the MLE reconstruction accuracy will be worse than the theoretical limit.
3.1.1 Validation with LUX experimental results
The LUX experiment operated a dual-phase xenon TPC to search for dark matter interactions between 2012 and 2016 [5]. The LUX TPC used 61 2" PMTs in the xenon gas phase and a mirrored array in the liquid to collect xenon scintillation and electroluminescence light. In this configuration, LUX reported a xyx-y position resolution of $\sim$2 cm for single electron pulses with $\sim$20 detected photons using the MLE method [13].
The LUX TPC was 50 cm in diameter with a 0.5 cm electroluminescence region between the liquid surface and the TPC anode, and the PMT array was placed 6 cm above the anode. Fig. 10 (left) compares the radial component of the LUX LRF [13] with our simulated LRF using the LUX TPC configuration  [^3] Therefore, this TPC model is different from our default 2” PMT TPC model used in other validations.. In addition to adjusting the TPC dimensions, we also increase the PMT center-to-center separation to 6.125 cm from the nominal 2" diameter to be consistent with LUX where the extra space was needed to support the PMTs. Despite our simplified optical model, the two LRF functions are nearly identical, which provides a strong validation of our optical simulation model.
Figure 10: Left - simulated light collection efficiency in this work in comparison with the radial component of the LUX LRF [13]; the best fit to our simulation using the LUX LRF formula is also shown. Right - predicted SS event reconstruction resolutions as a function of detected photons for 3 event positions (see text), plotted on top of experimental result from the LUX experiment. The LUX parameters are used in our simulation and sensitivity predictions.
Using the LUX radial LRF provided in [13], we carried out a similar theoretical study to predict the SS event reconstruction resolution in the LUX detector using FI. Here we only consider SS events near the center of the detector so the polar component of the LUX LRF (not included in this study) is small and can be neglected. Fig. 10 (right) shows the predicted SS reconstruction resolution curves for events right below the center PMT (red solid line), between 2 center PMTs (red dashed line) and at the center of 3 PMTs (red dotted line), which are referred to as positions A, B, and C, respectively in Fig. 3, on top of experimental results reported by LUX.
Because of the large light source-PMT array distance in the LUX configuration, the detector’s position resolution is not sensitive to the event position relative to the PMT layout, and thus the predictions for the 3 representative event positions are very similar. Notably, the FI predictions are only $\sim$20% better than those achieved by the LUX MLE reconstruction for actual data, strongly supporting the FI results as practical upper limits. The difference between FI predictions and LUX results is attributed to detector effects such as light reflection off detector walls, non-unit and non-uniform PMT quantum efficiencies, and additional PMT noise to Poissonian fluctuations. These experimental factors are not included in our theoretical studies. As a result, the simulation-based validations, which also ignore these effects, may produce better agreement with theory than for real experiments. However, the close match between simplified FI predictions and the LUX experimental results confirms that the theoretical precisions can be reasonably approached in real experiments.
3.2 Multi-site background rejection in the transverse plane
In this section, we seek to validate FI predictions for MS2 events using the PCA, MLE and CNN methods. Instead of estimating the inter-vertex separation for such events, we focus on the power of these methods in differentiating SS and MS2 events, which is the main goal of many physics applications. Similar to the SS reconstruction study, we simulate MS2 events by randomly drawing photons from two vertices of assigned xyx-y positions and estimating the PMT occupancy using the parameterized LRF. The PMT size is assumed to be 2" and the total number of photons are evenly split between the two vertices.
3.2.1 Principal component analysis
MS2 events are expected to produce a photon hit pattern with larger variance along the direction connecting the two vertices (hereafter referred to as the principal direction) compared to its orthogonal direction. We use the Principal Component Analysis (PCA) to identify the principal axis direction and to quantify this asymmetric variance for SS and MS events. Mathematically, PCA diagonalizes the covariance matrix and rotates the coordinates to align with the principal direction. In addition to computing the rotation angle, PCA also calculates the eigenvalues of the covariance matrix, which correspond to the variance in photon spread in the principal and secondary directions. We will refer to the square roots of these variance values as the principal and secondary spreads.
Figure 11 (left) shows the distributions of principal (blue) and secondary (black) spread for double vertices separated by 0, 3 cm and 6 cm with a total of 500 detected photons. Here we assume a sensor pixel size of 2" and an 1:1 photon splitting between the two vertices. Due to statistical fluctuations, the PCA algorithm proposes principal and secondary directions even for SS events (0 separation between vertices). Therefore, the appearance of asymmetric variance, by itself, cannot be used to discriminate SS and MS events. Instead, we use the displacement of the mean principal spread from the SS mean to estimate how well MS events may be distinguished from SS events. This method takes statistical fluctuations of SS variance distribution into consideration. When this displacement equals the width of the principal spread distribution, we obtain a 50% SS acceptance while allowing 16% MS contamination. The corresponding vertex separation can be directly compared to the benchmark MS2 resolution $\delta_{d}$ defined in Sec. 2.3. The PCA-derived $\delta_{d}$ values are compared to that predicted by FI, as a function of detected photon count, in Figure 11 (right). The PCA-derived resolution values follow the same trend as the theoretical predictions, but are approximately twice larger at the same photon statistics.
3.2.2 Likelihood ratio analysis
MLE can reconstruct the inter-vertex separation of MS2 events. However, when the total number of vertices is unknown as in a real MS event, the MLE optimization problem can become ill-defined. Therefore, we only quantify the goodness of MLE fits for simulated MS2 events with a likelihood ratio, which is defined as
$q=-2\textrm{log}\left(\frac{\mathcal{L}(\hat{m},\hat{x}_{1},\hat{y}_{1},f_{2}=0|n_{k})}{\mathcal{L}(\hat{m},\hat{x}_{1},\hat{y}_{1},\hat{f}_{2},\hat{d},\hat{\phi}|n_{k})}\right)$ (3.2)